OpenStudy (anonymous):

solve the system of equations 5x-y=27 5x+4=42

OpenStudy (anonymous):

Is there supposed to be a "y" in the second equation?

OpenStudy (anonymous):

ohh my bad its um 5x+4y=42

OpenStudy (anonymous):

Ok, what you can do is use the elimination method, since both equations have 5x, we can negate one of them. Lets negate the top equation and put it beneath the other one like this: 5x+4y=42 -5x+y=-27 ---------- Now we can add these equations going down and it'll look like this: 5y=15 y=3 Now that you have y, you can plug it into either equation to solve for x.

OpenStudy (anonymous):

5x+4(3)=42 5x+12=42 -12 -12 5x=30 x=6 (6,3)

OpenStudy (anonymous):

That looks good to be. You can check your answer by plugging the points into the original equations, or graphing them.

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

You're welcome, glad I could help